In 8 years time a boy will be 3 times the age he is now. How old is the boy now?
Found 3 solutions by math_helper, ankor@dixie-net.com, greenestamps:
Answer by math_helper(2461)

x+8 = 3x --> x = 4. The boy is 4yrs old now.

Answer by ankor@dixie-net.com(22740)

a = his present age
:
In 8 years time a boy will be 3 times the age he is now.
3a = a + 8
3a - a = 8
2a = 8
a = 8/2
a = 4 yrs old now

Answer by greenestamps(13200)



Here is a way to set up the problem for solving using algebra that is different and less common than the usual setup shown by the other two tutors. In certain similar problems, this alternative method for setting up the problem might lead to an easier solution.

Let x be the boy's current age; then his age when he is 3 times as old as he is now is 3x.

The difference between those two ages is 3x-x = 2x.

But the problem tells us the difference between those two ages is 8 years. So

2x = 8
x = 4

The boy's current age is x=4.